# An Adaptive Placement and Parallelism Framework for Accelerating RLHF Training

### v2

@@ Abstract
Recently, ChatGPT or InstructGPT like large language models (LLM) has made a significant impact in the AI world. Many works have attempted to reproduce the complex InstructGPT’s training pipeline, namely Reinforcement Learning with Human Feedback (RLHF). However, the mainstream distributed RLHF training methods typically adopt a fixed model placement strategy, referred to as the Flattening strategy. This strategy treats all four interdependent models involved in RLHF as a single entity, distributing them across all devices and applying parallelism techniques designed for a single model, regardless of the different workloads inherent to each model. As a result, this strategy exacerbates the generation bottlenecks in the RLHF training and degrades the overall training efficiency. To address these issues, we propose an adaptive model placement framework that offers two flexible model placement strategies. The Interleaving strategy helps reduce memory redundancy and communication costs of RLHF training by placing models without dependencies on exclusive devices with careful orchestration. On the other hand, the Separation strategy improves the throughput of model training by separating the training and inference runtime of the RLHF pipeline with additional shadow models. Furthermore, our framework provides a simple user interface and allows for the agile allocation of models across devices in a fine-grained manner for various training scenarios, involving models of varying sizes and devices of different scales. Extensive experiments have demonstrated that our Interleaving and Separation strategies can achieve notable improvements up to 11, compared to the current state-of-the-art (SOTA) approaches. The results highlight the effectiveness and adaptability of our approaches in accelerating the training of distributed RLHF.
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities across various natural language processing tasks[2, 5, 29, 35]. However, it is crucial to fine-tune LLM models to align them with human preferences. In the absence of proper alignment with human feedback, LLMs may exhibit behaviors that deviate from expectations. These behaviors include generating fabricated information, pursuing inaccurate objectives, and producing harmful, misleading, and biased expressions[18, 13, 8]). To tackle this issue, researchers have proposed a series of approaches aimed at ensuring that LLMs adhere to ethical considerations and societal norms.
One of the most successful approaches in addressing the alignment issue is Reinforcement Learning from Human Feedback [25, 1, 18]. Particularly, the pipeline proposed in InstructGPT [18], a sibling model of ChatGPT, has gained significant attention. RLHF combines human alignment with Reinforcement Learning (RL), a decision-making approach that allows AI agents to learn from interactions with the environment, receiving feedback in the form of rewards. This integration of RL with Human Feedback (HF) enables researchers to enhance the capabilities of LLMs and improve their alignment with human values. RLHF represents a significant advancement in building more reliable and responsible LLMs.
@@
The predominant RL algorithm to finetune LLM is Proximal Policy Optimization (PPO) [21]. It is the main process in the RLHF pipeline known as “Step-3: Training RL Policy” in InstructGPT. This process leverages an Actor-Critic architecture[16], comprising four interconnected models: a Reference (Ref) model, a Reward model, an Actor model, and a Critic model. The pipeline is segmented into three sequential stages—generation, forward, and training—as depicted in Figure 1. Note that Actor and Critic models are trainable models, which are leveraged for both training and inference runtime while Ref and Reward models are inference models utilized solely for inference. Specifically, in the generation stage, the Actor model generates the response from the query. In the subsequent forward stage, given the query and response, all four models follow a sequential execution order to produce intermediate output results such as logits or values. For example, the Actor, Reference, Reward, and Critic components conduct the forward stage one after the other. These results are stored in the experience replay buffer similar to DQN[17]. During the training stage, the Actor and Critic models are trained using batch data sampled from the buffer. This whole procedure repeats over a number of steps until converges. It is important to note that the Actor and Critic models utilize the same underlying model but under a mixed usage of training and inference runtime (inference are used in both generation or forward stage) across these three stages.
@@
Across the training of RLHF, the size of individual LM parameters can exhibit significant variations, e.g. spanning from 7 billion (7B) to 65 billion (65B) parameters in models like the Llama series [26]. The presence of LMs in RLHF pipeline amplifies memory usage quadratically, which underscores the critical need for efficient distributed training methodologies. However, traditional distributed training techniques, such as data parallelism or model parallelism[10, 32, 20, 24], are primarily designed for training a single model, not for efficiently managing multiple interdependent models as seen in RLHF. This necessitates the development of innovative model placement strategies.
@@
Existing works[33, 3, 27] typically employ a straightforward model placement strategy, namely Flattening strategy. This strategy treats all involved models as a single entity and applies standard parallelism techniques designed for individual models. While this strategy aligns well with the sequential execution order as the standalone PPO algorithm, it brings several limitations and opportunities for optimization. 1) This approach results in reduced memory available for a single model on each device. Therefore, we have to add up more devices and parallelism techniques to accommodate all four models, with increased memory redundancy and communication costs [20, 31]. 2) Different scales of model sizes require fine-grained model placement strategies. For instance, 7 billion-level models could be held by fewer devices compared with 65 billion-level models, which creates the opportunity for optimization. 3) The workloads during training and inference runtime inherently differ, necessitating distinct optimization techniques, such as distributed strategies. Specifically, in the training stage, it is advisable to shard the optimizer state, gradients, and parameters sequentially for better efficiency. On the other hand, during the inference stage, only the parameters need to be shared, as they are the only components utilized during inference.
@@
Therefore, we propose the Adpative Placement and Parallelism (APP) framework, which could significantly enhance the training efficiency in distributed RLHF training while ensuring no compromise on model performance. Firstly, our framework introduces two innovative model placement strategies to improve training efficiency in various RLHF training scenarios. Secondly, we design an APP Execution Engine abstraction that provides a user-friendly interface to modeling experts, shielding them from the intricate details of model placement and parallelism strategies.
@@
i) Adaptive Model Placement: Our APP framework introduces two flexible model placement strategies. Our Interleaving strategy places models without dependencies on exclusive devices and our Separation strategy decouples the training and inference runtime of the RLHF pipeline, significantly improving the training efficiency.
ii) GPU Heterogeneity Support: The Separation strategy accounts for mixed GPU capabilities, facilitating RLHF model training on a wider array of affordable and specialized GPU series for training or inference.
iii) Enhanced Performance: Our extensive experiments confirm that our method outperforms the state-of-the-art up to in terms of throughput.
iv) Ease of Use: The APP framework features an Execution Engine supporting various model placement strategies and acceleration techniques with minimal or no code changes. It makes the distributed RLHF training more accessible for users without expertise in distributed training.
Furthermore, we are actively working towards releasing the code for the APP framework as an open-source project.
@@ II-A Data and Model parallelism
According to different parallel objects, distributed training technology for parallelizing single language model can be divided into data parallelism and model parallelism.
Data Parallelism (DP) involves dividing the input data into multiple shards and assigning them to different devices for parallel processing. In traditional DP training methods like AllReduce[7, 22] or Distributed Data Parallel (DDP)[14, 19], each device holds a complete replica of the model states and uses the AllReduce primitive to synchronize the model states across devices. ZeRO (Zero Redundancy Optimizer)[20] is a technique that addresses memory redundancy of the model state replicas held on each device, which dominates the current LLM training. It categorizes the model states into three parts: optimizer states, gradients, and parameters. To optimize memory redundancy, ZeRO performs sequential optimization in three stages: ZeRO-1, ZeRO-2, and ZeRO-3. In ZeRO-1, only the optimizer states are sharded across devices, which requires communication in the parameter update stage. In ZeRO-2, the gradients are additionally sharded, introducing additional reduce-scatter communication after each backward stage. ZeRO-3 extendedly shards the model parameters to all devices, while it requires extra AllGather communication in both forward and backward stages. The increased level of ZeRO parallelism helps reduce memory redundancy but comes with increased communication costs[20, 31].
Model Parallelism, including Tensor and Pipeline Parallelism, involves distributing model parameters across multiple devices within or across instances[24]. Tensor Parallelism (TP)[23, 24] achieves this by vertically splitting the model, sharding tensors across multiple devices, and computing through distributed matrix multiplication. However, due to the frequent global communication, Tensor Parallelism is more efficient within a single node with a high GPU interconnect bandwidth. As the number of nodes increases, its scalability becomes limited. On the other hand, Pipeline Parallelism (PP)[10, 9] adopts a horizontal model splitting approach, assigning different layers to different devices. It usually uses micro-batching to hide the bubbles which is the idle duration of the device in the pipeline.
@@ II-B Heterogeneous Network
Clusters with heterogeneous networks are common in the majority of cloud computing environments[12], where the bandwidth between intra-node and inter-node is unbalanced. For instance, in cloud A100 GPU clusters, NVLink/NVSwitch with a bandwidth up to 4.8TGbps is configured within the node. However, the inter-node bandwidth, which might utilize whether InfiniBand or Ethernet, typically ranges from 200Gbps to 800Gbps. Such disparities mean that the inter-node bandwidth can be approximately 24 times slower than the intra-node bandwidth[31, 36, 28]. This limitation becomes even more pronounced in heterogeneous GPU clusters, such as a mixture of A100 and V100, where different types of GPU devices can only be connected using Ethernet. The significant gap in bandwidth within and between nodes greatly hampers the communication efficiency of Tensor Parallelism and ZeRO, especially at higher levels of ZeRO[31, 20]. ZeRO-3, for instance, requires frequent AllGather communication primitives across all nodes to collect model parameters which are shared across all devices, as compared to ZeRO-2. Meanwhile, tensor parallelism requires more global communications for each matrix computation during computation. As a result, the inter-intra-node bandwidth gap will significantly hamper the efficiency of parallelism techniques. Moreover, as the size of the cluster scales out, the communication overhead caused by heterogeneous networks grows larger.
III Related Works
Since the emergence of ChatGPT and InstructGPT, several distributed frameworks have been proposed to support the parallelization of complex InstructGPT-like RLHF training pipelines. These works include trlX [3], Colossalai [15], Deepspeed-Chat [33], and trl [27]. From an algothrim aspect, these works primarily consider two RLHF model structures: AC-Share[3] and AC-NonShare[33], depending on whether the Actor and Critic model share the parameters or not. For example, in the AC-Share structure, the Actor and Critic models share parameters by adding an extra linear layer to differentiate between them. This reduces the memory and resource requirements for RLHF training, but it may potentially lead to a loss in model performance.
To parallelize the training of multiple interdependent models in the RLHF pipeline, all of these works adopt a fixed model placement strategy or its variant, referred to as the Flattening strategy[33, 27, 15, 3]. The Flattening strategy straightforwardly takes four models as a whole and treats them equally. For instance, DeepSpeed-Chat[33], one of the most popular RLHF training frameworks using the AC-NonShare structure, places all four models of RLHF pipeline on each device and applies data and model parallelism techniques like ZeRO to parallelize the training. trlX[3], another well-known work utilizing the AC-Share structure, employs a variant of the Flattening placement strategy. It places the Reward model on a fully occupied device to leverage its computational capabilities exclusively while deploying the Actor, Critic (shared with the Actor model), and Ref model on other devices using the Flattening strategy.
@@ III Related Works
Overall, unlike standard approaches for training a single model, distributed RLHF training involves managing the model placement and computation of multiple models across both training and inference runtimes. This gap creates an opportunity for further optimization to improve the training efficiency of RLHF training.
III-A Limitations in Existing Distributed RLHF Training
@@
Without loss of generality, we take the AC-Share structure as an example. Figure 1 illustrates that the Actor Generation accounts for more than 85% of the total duration in one step, while the Training stage only occupies 10% and then other forward stages using DeepSpeed-Chat. Clearly, the generation stage significantly slows down the entire RLHF training. The bottleneck comes from two main aspects: sequential execution and the fixed model placement strategy.
@@
Sequential Execution. Existing works typically follow a sequential execution order, similar to the single-machine RLHF algorithm as shown in Figure 1, where each stage or model is executed one after another. However, the generation stage is essentially more computation-intensive than other stages. The generation stage primarily involves a time-consuming auto-regressive generation using the Actor model, followed by four sequential forward phases in each model. In this sequential execution order, the intermediate results generated from the time-cost generation stage in the experience replay buffer are rapidly consumed by the following stages, namely the forward and training stages. However, the training model consumes typically eight times memory as much as the inference model. Consequently, the training models are mostly idle but occupy a significant amount of memory.
@@
Fixed Model Placement Strategy. These works primarily rely on a fixed model placement strategy, namely the Flattening strategy, regardless of the size of the models used. It is easy to implement without additional intermedidate data interaction and ensures that all devices keep working without idleness throughout all three stages but depends on sequential execution 111Someone may argue another simple placement strategy to remove this “apparent” sequential execution order. It could exclusively place four models engaged in the Forward Stage on separate groups of devices for parallelization, given non-data dependencies among these four models during the forward stage. Nonetheless, this strategy fails to consider the interdependencies between the models across three stages. For example, the Actor model will be also utilized in the generation and training stage. During the generation stage, only devices with Actor models work, leaving other devices idle. The Flattening Strategy is more efficient than this strategy by fully utilizing all devices but heavily relies on sequential execution. However, this lack of adaptability does possess three limitations: 1) The Flattening strategy incurs memory redundancy and additional communication costs in RLHF training. It deploys all four models on every device, significantly reducing the available memory on each device for a single model. This necessitates an increased level of ZeRO parallelism to accommodate all models but with extra memory redudancy and communication costs, as discussed in section II-A. 2) This strategy fails to decouple the training and inference runtime, which naturally differ from each other and require distinct optimization techniques. In terms of parallelism, the training phase requires taking all the model states including parameters, gradients, and optimizer states into consideration, while only parameters are used in the inference phase, which requires different parallelism techniques. For example, Tensor Parallelism or PaRO[24, 31] could be used to parallelize the model inference rather than ZeRO parallelism which fits model training. 3) Heterogeneous GPU support is difficult under this strategy. It is due to kernel compatibility and network heterogeneity issues[6], hindering the usage of more affordable and specialized GPU series. Also, obtaining heterogeneous GPUs (e.g., a combination of A100 and V100) is comparatively easier than scheduling hundreds of homogeneous high-end GPUs in the industry[11, 30].
III-B Key Challenges

### v3

@@ Abstract
Recently, ChatGPT or InstructGPT like large language models (LLM) has made a significant impact in the AI world. Many works have attempted to reproduce the complex InstructGPT’s training pipeline, namely Reinforcement Learning with Human Feedback (RLHF). However, the mainstream distributed RLHF training methods typically adopt a fixed model placement strategy, referred to as the Co-located strategy. This strategy treats all four interdependent models involved in RLHF as a single entity, distributing them across all devices and applying parallelism techniques designed for a single model, regardless of the workload heterogeneity inherent to each model. As a result, this strategy exacerbates the generation bottlenecks in the RLHF training and degrades the overall training efficiency. To address these issues, we propose a flexible model placement framework that offers two general and agile model placement strategies. The Interleaving strategy helps reduce memory redundancy and communication costs of RLHF training by placing models without dependencies on exclusive devices with careful orchestration. On the other hand, the Disaggregated strategy improves the throughput of model training by separating the training and inference runtime of the RLHF pipeline with additional shadow models. Furthermore, our framework provides a simple user interface and guidelines to easily and flexibly configure these strategies in various training scenarios. Our experiments have shown that our strategy can achieve notable improvements up to 11, compared to the current state-of-the-art (SOTA) approaches. The results highlight the effectiveness and adaptability of our methods in accelerating the training of distributed RLHF.
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities across various natural language processing tasks [4, 15, 42, 5, 35]. However, it is crucial to fine-tune LLM models to align them with human preferences. In the absence of proper alignment with human feedback, LLMs may exhibit behaviors that deviate from expectations. These behaviors include generating fabricated information, pursuing inaccurate objectives, and producing harmful, misleading, and biased expressions [24, 18, 10, 38]). To tackle this issue, researchers have proposed a series of approaches aimed at ensuring that LLMs adhere to ethical considerations and societal norms. One of the most successful approaches in addressing the alignment issue is Reinforcement Learning from Human Feedback (RLHF) [33, 1, 3, 6]. Particularly, the pipeline proposed in InstructGPT [24], a sibling model of ChatGPT, has gained significant attention. RLHF combines human alignment with Reinforcement Learning (RL) [34, 13, 23], a decision-making approach that allows AI agents to learn from interactions with the environment, receiving feedback in the form of rewards. This integration of RL with Human Feedback (HF) enables researchers to enhance the capabilities of LLMs and improve their alignment with human values.
@@
The predominant RL algorithm to finetune LLM is Proximal Policy Optimization (PPO) [28, 24, 26] and is known as “Step-3: Training RL Policy” in InstructGPT. This process leverages an Actor-Critic architecture [22, 34, 24], comprising four interconnected models: a Reference (Ref) model, a Reward model, an Actor model, and a Critic model. The pipeline is segmented into three sequential stages—generation, forward, and training—as depicted in Fig. 1. Note that Actor and Critic models are trainable models, which are leveraged for both training and inference runtime while Ref and Reward models are inference models utilized solely for inference. Specifically, in the generation stage, the Actor model generates the response from the query. In the subsequent forward stage, given the query and response, all four models follow a sequential execution order to produce intermediate output results such as logits or values. For example, the Actor, Reference, Reward, and Critic models conduct the forward stage one after the other. These results are stored in the experience replay buffer. During the training stage, the Actor and Critic models are trained using batch data sampled from the buffer. This whole procedure repeats over a number of steps until it converges. It is important to note that the Actor and Critic models utilize the same underlying model but under a mixed usage of training and inference runtime across these three stages.
@@
Across the training of RLHF, the size of individual LM parameters can exhibit significant variations, e.g. spanning from 7 billion (7B) to 65 billion (65B) parameters in models [35, 8]. The presence of LMs in the RLHF pipeline amplifies memory usage quadratically, which underscores the critical need for efficient distributed training methods. However, distributed training techniques, such as data parallelism or model parallelism [41, 27, 31], are primarily designed for training a single model, not for efficiently managing multiple interdependent models as seen in RLHF. This necessitates the development of innovative model placement strategies.
@@
Existing works [43, 12, 21] typically employ a straightforward model placement strategy, namely Co-located strategy. This strategy treats all involved models as a single entity and applies standard parallelism techniques designed for individual models. While this strategy aligns well with the sequential execution order as the single-machine PPO algorithm, it brings several limitations and opportunities for optimization. 1) This approach results in reduced memory available for a single model on each device. Therefore, we have to add up more devices and parallelism techniques to accommodate all four models, with increased memory redundancy and communication costs [27, 40, 32]. 2) Different scales of model sizes require fine-grained model placement strategies. For instance, 7B models could be held by fewer devices compared with 65B models, which creates the opportunity for optimization. 3) The workloads during training and inference runtime inherently differ, necessitating distinct optimization techniques [40, 19]. Specifically, in the training stage, it is advisable to shard the optimizer state, gradients, and parameters sequentially for better efficiency. However, only the parameters need to be shared during the inference stage, as gradients and optimizer state are not required during inference [2].
@@
Therefore we propose FlexRLHF, a flexible Reinforcement Learning from Human Feedback (RLHF) framework, that significantly enhances the training efficiency in distributed RLHF training while ensuring no compromise on model performance. Firstly, our framework introduces two innovative model placement strategies to improve training efficiency in various RLHF training scenarios. Secondly, we design a FlexRLHF Execution Engine abstraction that provides a user-friendly interface to modeling experts, shielding them from the intricate details of model placement and parallelism strategies. Thirdly, we provide a guideline on how to efficiently configure these distributed strategies in different training scenarios.
@@
i) Flexible Model Placement: Our FlexRLHF framework introduces two flexible model placement strategies. Our Interleaving strategy places models without dependencies on exclusive devices and our Disaggregated strategy decouples the training and inference runtime of the RLHF pipeline, significantly improving the training efficiency.
ii) Heterogeneity Support: Our framework accommodates diverse runtime environments, including different distributed training engines (e.g., DeepSpeed and Megatron) and inference engines. It also supports GPUs with varied capabilities and costs.
iii) Superior Performance: Our extensive experiments confirm that our method outperforms the state-of-the-art up to in terms of throughput.
iv) Ease of Use: Our framework features an Execution Engine with guidelines to support various model placement strategies and acceleration techniques with minimal or no code changes.
@@ II-A Data and Model parallelism
According to different parallel objects, distributed training technology for parallelizing single language models can be divided into data parallelism and model parallelism.
Data Parallelism (DP) involves dividing the input data into multiple partitions and assigning them to different devices for parallel processing. In traditional DP training methods like AllReduce [9, 29] or Distributed Data Parallel (DDP) [20, 25], each device holds a complete replica of the model states and performs the AllReduce primitive to synchronize the model states across devices. ZeRO (Zero Redundancy Optimizer) [27] is a technique that addresses memory redundancy of the model state replicas held on each device, which dominates the DP-based LLM training. However, the increased level of ZeRO parallelism helps reduce memory redundancy but comes with growing communication costs [27, 40].
Model Parallelism, including Tensor and Pipeline Parallelism, involves distributing model parameters across multiple devices within or across instances [31]. Tensor Parallelism (TP) [30, 31] achieves this by vertically splitting the model, sharding tensors across multiple devices, and computing through distributed matrix multiplication. However, due to the frequent global communication, Tensor Parallelism is more efficient within nodes with a high GPU interconnect bandwidth. On the other hand, Pipeline Parallelism (PP) [14, 11] adopts a horizontal model splitting approach, assigning different layers to different devices. It usually uses micro-batching to hide the bubbles which is the idle duration of the device in the pipeline.
@@ II-B Heterogeneous Network
Clusters with heterogeneous networks are common in cloud computing environments [17], where the bandwidth between intra-node and inter-node is unbalanced. Typically, the inter-node bandwidth can be from 3x to 24x times slower than the intra-node bandwidth [44, 37]. This limitation becomes even more pronounced in heterogeneous GPU clusters, such as a mixture of A100 and V100, where different types of GPU devices can only be connected using Ethernet. The significant gap in the network between intra- and inter-node greatly hampers the communication efficiency of Tensor Parallelism and Data Parallelism, especially at higher levels of ZeRO [40, 27]. Moreover, as the size of the cluster scales out, the communication overhead caused by heterogeneous networks grows larger, which downgrades the training efficiency.
III Related Works and Motivation
III-A Related Works
Since the emergence of ChatGPT and InstructGPT, several distributed frameworks have been proposed to support the parallelization of complex InstructGPT-like RLHF training pipelines [43, 36, 21, 12]. From a modeling aspect, these works primarily consider two RLHF model structures: AC-Share [12] and AC-NonShare [43], depending on whether the Actor and Critic model share the parameters or not. For example, in the AC-Share structure, the Actor and Critic models share parameters by adding an extra linear layer to differentiate between them, and vice versa. This reduces the memory and resource requirements for RLHF training, but it may potentially lead to a loss in model performance.
To parallelize the training of multiple interdependent models in the RLHF pipeline, all of these works adopt a fixed model placement strategy or its variant, referred to as the Co-located strategy [43, 36, 21, 12]. The Co-located strategy straightforwardly takes four models as a whole and treats them equally. For instance, DeepSpeed-Chat [43], one of the most popular RLHF training frameworks using the AC-NonShare structure, places all four models of RLHF pipeline on each device and applies data and model parallelism techniques like ZeRO to parallelize the training. trlX [12], another well-known work utilizing the AC-Share structure, employs a variant of the Co-located placement strategy. It places the Reward model on a fully occupied device to leverage its computational capabilities exclusively while deploying the Actor, Critic (shared with the Actor model), and Ref model on other devices using the Co-located strategy.
@@ III-A Related Works
In addition, Deepspeed-Chat further proposes a Hybrid Engine [43]. During the training phase, it utilizes ZeRO for efficient memory management, while switching to tensor parallelism for the generation phase. To reshard model states between training and inference runtime, it performs the AllGather operation to gather the whole parameters into tensor parallel workers in each generation. Although it has avoided frequent AllGather communication in each forward for one generation, the remaining one for each generation still introduces significant overhead.
III-B Limitations in Existing Distributed RLHF Training
@@
Without loss of generality, we take the AC-Share structure as an example. Fig. 1 illustrates that the Actor Generation accounts for more than 85% of the total duration in one step, while the Training stage only occupies 10% and then other forward stages when using DeepSpeed-Chat. The generation stage significantly slows down the entire RLHF training. The bottleneck comes from two main aspects: sequential execution and the fixed model placement strategy.
@@
Sequential Execution. Existing works typically follow a sequential execution order, similar to the single-machine RLHF algorithm as shown in Fig. 1, where each stage or model is executed one after another. However, the generation stage is essentially more computation-intensive than other stages. The generation stage primarily involves a time-consuming auto-regressive generation using the Actor model, followed by four sequential forward phases in each model. In this sequential execution order, the intermediate results generated from the time-cost generation stage in the experience replay buffer are rapidly consumed by the following stages, namely the forward and training stages. However, the training model consumes typically eight times memory as much as the inference model. Consequently, the training models are mostly idle but occupy a large amount of memory.
@@
Fixed Model Placement Strategy. These works primarily rely on a fixed model placement strategy, namely the Co-located strategy, regardless of the size of the models used. It is easy to implement without additional intermediate data interaction and ensures that all devices keep working throughout all three stages but depends on sequential execution. 111Someone may argue a simple placement strategy to remove the sequential execution order. It could exclusively place four models engaged in the Forward stage on separate groups of devices for parallelization, given non-data dependencies among these four models during the forward stage. Nonetheless, this strategy fails to consider the interdependencies between the models across three stages. For example, the Actor model will be also utilized in the generation and training stage. During the generation stage, only devices with Actor models work, leaving other devices idle. The Co-located Strategy is more efficient than this strategy, as it ensures all devices are actively engaged across all three stages. However, as discussed, it results in additional memory redundancy and costly inter-node communication, which downgrades the throughput. However, this lack of adaptability does possess three limitations:
1) The Co-located strategy incurs memory redundancy and additional communication costs in RLHF training. It deploys all four models on every device, significantly reducing the available memory on each device for a single model. This necessitates an increased level of ZeRO parallelism to accommodate all models but with extra memory redundancy and communication costs, as discussed in section II-A. 2) This strategy fails to decouple the training and inference runtime, which naturally differ from each other and require distinct optimization techniques. Moreover, from a parallelism perspective, the training phase necessitates consideration of all model states, whereas the inference phase utilizes only parameters, thus demanding different parallelism approaches. In contrast to a hybrid approach, we argue that it is inefficient to combine different techniques in a hybrid engine. This inefficiency stems not only from the aforementioned communication challenges but also from the inability to leverage more effective, specialized optimization techniques for distinct runtimes. For instance, Megatron could be employed for training, while vLLM could be utilized for generation [19]. 3) Heterogeneous GPU support is difficult under this strategy. It is due to kernel compatibility and network heterogeneity issues [7], hindering the usage of more affordable and specialized GPU series. Also, obtaining heterogeneous GPUs (e.g., a mixture of A100 and V100) is comparatively easier than homogeneous high-end GPUs in the cluster [16, 39].
III-C Key Challenges